State whether the following statements are true or false. Give reasons for your answers.
a. the function f:R³→R, givan by f(x,y,z)=|x|+|y|+|z| is differential at (2,3,-1)
1
Expert's answer
2020-05-19T08:54:38-0400

Consider "f" on "S=(0,+\infty)\times(0,+\infty)\times(-\infty,0)".

We have "f(x,y,z)=x+y-z" on "S".

So "f" is differentiable on "S". Since "S" is an open set, and "(2,3,-1)\in S", we obtain that "f" is differentiable at "(2,3,-1)"

Answer: "f" is differentiable at "(2,3,-1)"